The sum of two numbers is 141. one number is 3 more than five times the other. find the two number.
Answer by deborabr(181)

x+y=141
x=3+5y
substituting x in the first equation
3+5y+y=141
6y=141-3
6y=138
y=138/6
y=23
now substituting y in the second equation to find x
x=3+5*23
x=3+115
x=118
the two numbers are 23 and 118